OpenStudy (anonymous):

1. Solve the system of equations below algebraically. 2x + 3y = 6 -5x + 2y = 4

OpenStudy (anonymous):

2) 5y + 3x = -10 \[-\frac{ 5 }{ 3 }y = x - \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

You need to solve for x as it relates to y in one of the equations. Then plug that value into the second equation. Thus if 2x + 3y= 6, then 2x = -3y +6 or x= -3/2y +3. Take this and plug it into the second equation to solve for y. -5 (-3/2y +3) +2y =4. Once you solve for y you can get x.

OpenStudy (anonymous):

2x + 3y = 6 -5x + 2y = 4 10x + 15y = 30 -10x + 4y = 8 _________________ 0x +19y = 38 y = 2 2x + 3 (2) = 6 2x = 6 -6 x = 0